At 0.0°C and 1.00 atm, 1.00 mol of a gas occupies a volume of 0.0224 m3.

 What is the number density?


1
Expert's answer
2022-05-18T15:28:18-0400
"P=nkT"

"n=\frac{P}{kT}=\frac{101.3*10^3}{1.38*10^{-23}*273}=2.7*10^{25}\:\rm m^{-3}"